# Triple DES AES Advanced Encryption Standard

### v2

@@
Qifan Wang University of BirminghamBirminghamUK q.wang.3@bham.ac.uk  and  David Oswald University of BirminghamBirminghamUK d.f.oswald@bham.ac.uk
@@ Abstract.
ccs: General and reference Surveys and overviewsccs: Computer systems organization Heterogeneous (hybrid) systemsccs: Security and privacy Hardware attacks and countermeasuresccs: Security and privacy Trusted computing
@@ 1. Introduction
To ensure the security and confidentiality of heterogeneous computing systems, researchers have focused on adapting privacy-preserving techniques traditionally used on CPUs to accelerators like GPUs. These techniques include cryptographic methods such as Homomorphic Encryption (HE(Mishra et al., 2020; Ng and Chow, 2021; Fan et al., 2023) and Secure Multi-Party Computation (MPC(Mishra et al., 2020; Ng and Chow, 2021; Tan et al., 2021; Watson et al., 2022; Jawalkar et al., 2023), as well as TEEs tailored for accelerators like GPUs (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; NVIDIA, 2022a; Jiang et al., 2022; Lee et al., 2022; Yudha et al., 2022; Deng et al., 2022; Ivanov et al., 2023; Mai et al., 2023; Wang et al., 2024b). TEEs offer certain advantages over cryptographic methods, particularly in performance and ease of deployment, providing a robust hardware-based security mechanism that is relatively straightforward to implement and manage. However, in contrast to cryptographic techniques, the security of TEEs relies on trust in the manufacturer and is difficult to analyse and prove rigorously.
@@
Recent research has explored extending TEEs to accelerators like GPUs, both for x86 (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; NVIDIA, 2022a; Lee et al., 2022; Yudha et al., 2022; Ivanov et al., 2023; Mai et al., 2023) and Arm (Jiang et al., 2022; Deng et al., 2022; Wang et al., 2024b) architectures. Most of these approaches involve modifying existing CPU TEEs (especially process-based TEEs like Intel Software Guard Extensions (SGX) and Arm TrustZone (TrustZone)) to secure the GPU software stack, such as GPU drivers, and proposing strategies to ensure the isolation and confidentiality of GPU kernel executions. Apart from this, NVIDIA has introduced a pioneering commercial platform in its Hopper architecture (NVIDIA, 2022a), which supports confidential computing through features like memory encryption and isolated execution contexts. However, the security threats associated with extending TEEs across both CPUs and accelerators in heterogeneous computing systems remain uncertain and require further investigation.
@@
Figure 1. Current state of security research on heterogeneous computing systems with GPU. GPU Crypto refers to papers that protect the privacy of GPU-based applications using cryptographic techniques such as HE and MPC. AA, MSCA, MCCA, PSCA, SFIA refer to architectural, microarchitectural side-channel/covert-channel, physical side-channel, and software-based fault injection attacks.
@@
1999200420092014201920242029 First GPU, NVIDIA GeForce256 ARM TrustZone Launched in ARM11 First Virtualized GPU (NVIDIA VGX) First Attack Exploiting GPU Vuln. (Lee et al., 2014; Maurice et al., 2014) First MSCA on GPU (Luo et al., 2015) Intel SGX Launched in Skylake First MCCA on GPU (Naghibijouybari et al., 2017) AMD SEV Launched in EPYC First FIA on GPU (Frigo et al., 2018) First x86-based GPU TEE (Volos et al., 2018) First MIG-enabled GPU, NVIDIA A100 Intel TDX Announced AMD SEV-SNP Available in Milan First Arm-based GPU TEE (Deng et al., 2022) First Commercial GPU TEE (NVIDIA, 2022a) First MCCA on MIG-enabled GPU (Zhang et al., 2023) First Arm CCA-based GPU TEE (Wang et al., 2024b)
@@ 2.1. General Purpose Graphics Processing Units
Accelerator Year Features Virtualization Memory Sharing Memory Size
None L1/L2 Cache DRAM L1 L2 DRAM
@@
Integrated/Discrete GPU. Integrated GPUs, such as Intel HD/Iris/UHD/Xe Graphics, Arm, and Qualcomm GPUs listed in Table 1, are integrated on the same die or package as the CPU. These GPUs share the system’s memory with the CPU. Discrete GPUs are standalone and dedicated graphics processing units separate from the CPU. They have their dedicated memory, which is distinct from the system memory (e.g., NVIDIA and AMD GPUs listed in Table 1).
@@ 2.2. Types of GPU Attacks
2.2.1. Architectural/Logical Attacks
These attacks encompass a class of attacks that exploit vulnerabilities or weaknesses in the underlying hardware architecture. These attacks aim to compromise system by targeting the physical or logical design of key hardware elements, including processing units (e.g., processors, hardware accelerators, and co-processors), the memory subsystem, and communication components responsible for interfacing with other system elements (Ghasempouri et al., 2023).
@@ 2.3.3. Virtualization-based TEEs
Arm CCA. Arm CCA is an advanced isolation technology introduced in Armv9. It preserves TrustZone’s NW and SW while introducing a new realm world designed for running multiple confidential “realms”. These realms, though isolated from other worlds, are managed by untrusted software components (e.g., hypervisor). A lightweight Realm Management Monitor (RMM(Arm, 2023b), operating within the hypervisor layer of the realm world, ensures memory isolation between realms. CCA incorporates a root world to accommodate the highest privilege monitor code stored in firmware (Wang et al., 2024b).
@@ 3. Overview and Taxonomy
forked edges, for tree= grow=east, reversed=true,anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hiddendraw, rounded corners,align=left, minimum width=2.5em, s sep=3pt, inner xsep=2pt, inner ysep=1pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=5.8em,font=,, where level=2text width=5.6em,font=, where level=3text width=9.0em,font=,[Confidential Computing on Heterogeneous CPU-GPU Systems, ver [GPU TEEs
(Section 4) [x86-based
GPU TEEs [Graviton (Volos et al., 2018), HIX (Jang et al., 2019), HETEE (Zhu et al., 2020), NVIDIA H100 (NVIDIA, 2022a), TNPU (Lee et al., 2022), LITE (Yudha et al., 2022), SAGE (Ivanov et al., 2023), Honeycomb (Mai et al., 2023),leaf,text width=38em] ] [Arm-based
GPU TEEs [StrongBox (Deng et al., 2022), CRONUS (Jiang et al., 2022), CAGE (Wang et al., 2024b),leaf,text width=15em] ] ] [Security of GPUs
(Section 5) [Architectural/
Logical Attacks [Lee et al. (Lee et al., 2014), Maurice et al. (Maurice et al., 2014), Pietro et al. (Pietro et al., 2016), Zhou et al. (Zhou et al., 2016), HE et al. (HE et al., 2020),leaf,text width=30em] ] [Microarchitectural
Side/Covert-
Channel Attacks [Attacks on Integrated GPU [Frigo et al. (Frigo et al., 2018), HE et al. (HE et al., 2020), Dutta et al. (Dutta et al., 2021), Taneja et al. (Taneja et al., 2023), Wang et al. (Wang et al., 2024a),leaf,text width=31em] ] [Attacks on Discrete GPU [Jiang et al. (Jiang et al., 2016), Jiang et al. (Jiang et al., 2017), Ahn et al. (Ahn et al., 2021a), Giner et al. (Giner et al., 2024), Luo et al. (Luo et al., 2019), Zhu et al. (Zhu et al., 2021),
Hu et al. (Hu et al., 2020), Naghibijouybari et al. (Naghibijouybari et al., 2017), Wei et al. (Wei et al., 2020), Ahn et al. (Ahn et al., 2021b), Nayak et al. (Nayak et al., 2021),
Cronin et al. (Cronin et al., 2021), Yang et al. (Yang et al., 2022), Dutta et al. (Dutta et al., 2023), Zhang et al. (Zhang et al., 2023),leaf,text width=32em] ] [Countermeasures [Wang and Lee (Wang and Lee, 2007), Kong et al. (Kong et al., 2009), Liu et al. (Liu et al., 2016), Zhang et al. (Zhang et al., 2023), Wang et al. (Wang et al., 2014),
Shafiee et al. (Shafiee et al., 2015), Di et al. (Di et al., 2021), Kadam et al. (Kadam et al., 2018), Kadam et al. (Kadam et al., 2020), Lin et al. (Lin et al., 2019),leaf,text width=30em] ] ] [Physical Side-
Channel Attacks [Physical Access-based [Luo et al. (Luo et al., 2015), Gao et al. (Gao et al., 2018), Hu et al. (Hu et al., 2020), Chmielewski and Weissbart (Chmielewski and Weissbart, 2021), Maia et al. (Maia et al., 2022),
Horvath et al. (Horvath et al., 2023), Taneja et al. (Taneja et al., 2023),leaf,text width=32em] ] [Remote Access-based [Zhan et al. (Zhan et al., 2022), Liang et al. (Liang et al., 2022),leaf,text width=12em] ] [Countermeasures [Mangard et al. (Mangard et al., 2008), Zhan et al. (Zhan et al., 2022),leaf,text width=13em] ] ] [Fault Injection
Attacks [Software-based [Frigo et al. (Frigo et al., 2018), Sabbagh et al. (Sabbagh et al., 2020), Sabbagh et al. (Sabbagh et al., 2021), Sun et al. (Sun et al., 2023), Santos et al. (Santos et al., 2021),leaf,text width=32em] ] ] ] [Security of CPU
TEEs (Section 6) [Security of Intel
SGX [Controlled Channel/Address
Translation-based Attacks [Xu et al. (Xu et al., 2015), Shinde et al. (Shinde et al., 2016), Van Bulck et al. (Van Bulck et al., 2017a), Van Bulck et al. (Van Bulck et al., 2017b), Wang et al. (Wang et al., 2017),
Gyselinck et al. (Gyselinck et al., 2018), Kim et al. (Kim et al., 2019), Fei et al. (Fei et al., 2021),leaf,text width=33em] ] [Cache-based Attacks [Götzfried et al. (Götzfried et al., 2017), Moghimi et al. (Moghimi et al., 2017), Schwarz et al. (Schwarz et al., 2017), Brasser et al. (Brasser et al., 2017b), Dall et al. (Dall et al., 2018),
Moghimi et al. (Moghimi et al., 2019),leaf,text width=32em] ] [Branch Prediction Attacks [Fei et al. (Fei et al., 2021), Lee et al. (Lee et al., 2017), Chen et al. (Chen et al., 2019), Evtyushkin et al. (Evtyushkin et al., 2018), Huo et al. (Huo et al., 2020), Koruyeh et al. (Koruyeh et al., 2018),leaf,text width=34em] ] [Transient Execution Attacks [Chen et al. (Chen et al., 2019), Van Bulck et al. (Van Bulck et al., 2018), Van Schaik et al. (Van Schaik et al., 2020), Ragab et al. (Ragab et al., 2021),leaf,text width=27em] ] [Software-based Fault
Injection Attacks [Murdock et al. (Murdock et al., 2020), Kenjar et al. (Kenjar et al., 2020), Qiu et al. (Qiu et al., 2020), Jang et al. (Jang et al., 2017),leaf,text width=24em] ] [Hardware-based Attacks [Chen et al. (Chen et al., 2021), Chen and Oswald (Chen and Oswald, 2023), Lee et al. (Lee et al., 2020),leaf,text width=18em] ] [Interface-based Attacks [Van Bulck et al. (Van Bulck et al., 2019), Alder et al. (Alder et al., 2020), Cloosters et al. (Cloosters et al., 2020), Khandaker et al. (Khandaker et al., 2020), Alder et al. (Alder et al., 2024),leaf,text width=32em] ] ] [Security of
Arm TrustZone [Implementation Bugs [Beniamini (Beniamini, 2016), Shen (Shen, 2015), Komaromy (Komaromy, 2018),leaf,text width=15em] ] [Cache-based Attacks [Zhang et al. (Zhang et al., 2016c), Guanciale et al. (Guanciale et al., 2016), Lipp et al. (Lipp et al., 2016), Zhang et al. (Zhang et al., 2016b), Cho et al. (Cho et al., 2018),leaf,text width=30em] ] [Branch Prediction Attacks [Ryan (Ryan, 2019),leaf,text width=4em] ] [Software-based Fault
Injection Attacks [Tang et al. (Tang et al., 2017), Qiu et al. (Qiu et al., 2020), Carru (Carru, 2017),leaf,text width=16em] ] ] [Security of
AMD SEV [Memory Encryption Issues [Hetzelt and Buhren (Hetzelt and Buhren, 2017), Buhren et al. (Buhren et al., 2017), Du et al. (Du et al., 2017), Morbitzer et al. (Morbitzer et al., 2018), Li et al. (Li et al., 2021b),
Li (Li, 2022), Li et al. (Li et al., 2021a), Li et al. (Li et al., 2021c),leaf,text width=32em] ] [Software-based Fault
Injection Attacks [Morbitzer et al. (Morbitzer et al., 2021), Buhren et al. (Buhren et al., 2021), Zhang et al. (Zhang et al., 2024), Schlüter et al. (Schlüter et al., 2024a), Schlüter et al. (Schlüter et al., 2024b),leaf,text width=33em] ] [Interrupt-driven Attacks [Wilke et al. (Wilke et al., 2024),leaf,text width=6em] ] ] ] [Potential Attacks
against GPU TEEs
(Section 7) [x86-based
GPU TEE [Existing Attacks against Intel SGX,leaf,text width=12em] [ Dynamic Voltage and Frequency Scaling (DVFS)-based Attacks,leaf,text width=21em] [Physical Attacks on PCIe bus and GPU Memory,leaf,text width=17em] [Architectural/Logical attacks on GPU,leaf,text width=13em] ] [NVIDIA Hopper
H100-based
GPU TEE [Existing attacks against VM-based TEEs,leaf,text width=14em] [DVFS-based Attacks,leaf,text width=8em] [Microarchitectural Side/Covert-Channel Attacks on GPU,leaf,text width=19em] ] [Arm-based
GPU TEE [Existing attacks against Arm TrustZone,leaf,text width=13em] [Microarchitectural Side/Covert-Channel Attacks on Integrated GPU,leaf,text width=22em] [Physical Attacks on GPU Memory,leaf,text width=12em] [DVFS-based Attacks,leaf,text width=8em] ] ] ]
Figure 3. Taxonomy of confidential computing on heterogeneous CPU-GPU systems
@@
In Figure 3, we summarise our taxonomy underlying this survey based on relevant related work. In the remainder of this paper, we first review and discuss existing GPU TEEs in Section 4, covering their threat models, hardware modifications, TCB size, and performance overhead. In Section 5, we give an overview of existing attacks targeting GPUs based on the attack types listed in Section 2. We also review state-of-the-art attacks against existing CPU TEEs, such as Intel SGX, Arm TrustZone, and AMD SEV in Section 6. Finally, based on these GPU TEE designs and the summarized attacks, we discuss potential attacks against existing (and future) GPU TEEs in Section 7.
@@ 4.1. Trusted Execution Environments on GPUs
[…]
@@
[…]
@@
[…]
@@
To support confidential GPU computing on next-generation Arm devices, CAGE (Wang et al., 2024b) presents a new scheme for Arm CCA’s realm-style architecture. It offloads resource-intensive but data-independent tasks to NW, while protecting sensitive data within realms through a shadow task mechanism. By configuring Granule Protection Check (GPC), CAGE restricts access from untrusted components. Additionally, it isolates realms in GPU computing by providing each realm with a distinct GPU memory view in the GPU GPC. CAGE retains high hardware compatibility, as it requires no additional hardware or modifications to existing CPU and GPU configurations.

### v3

@@
Qifan Wang Durham UniversityDurhamUnited Kingdom University of BirminghamBirminghamUnited Kingdom qifan.wang2@durham.ac.uk and David Oswald Durham UniversityDurhamUnited Kingdom University of BirminghamBirminghamUnited Kingdom david.f.oswald@durham.ac.uk
@@ Abstract.
journal: CSURjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1journalyear: 2026copyright: acmlicensedccs: General and reference Surveys and overviewsccs: Computer systems organization Heterogeneous (hybrid) systemsccs: Security and privacy Hardware attacks and countermeasuresccs: Security and privacy Trusted computing
@@ 1. Introduction
To ensure the security and confidentiality of heterogeneous computing systems, researchers have focused on adapting privacy-preserving techniques traditionally used on CPUs to accelerators like GPUs. These techniques include cryptographic methods such as Homomorphic Encryption (HE(Mishra et al., 2020; Ng and Chow, 2021; Fan et al., 2023) and Secure Multi-Party Computation (MPC(Mishra et al., 2020; Ng and Chow, 2021; Tan et al., 2021; Watson et al., 2022; Jawalkar et al., 2024), as well as TEEs tailored for accelerators like GPUs (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; NVIDIA, 2022a; Jiang et al., 2022; Lee et al., 2022; Yudha et al., 2022; Deng et al., 2022; Ivanov et al., 2023; Mai et al., 2023; Wang et al., 2024a). TEEs offer certain advantages over cryptographic methods, particularly in performance and ease of deployment, providing a robust hardware-based security mechanism that is relatively straightforward to implement and manage. However, in contrast to cryptographic techniques, the security of TEEs relies on trust in the manufacturer and is difficult to analyse and prove rigorously.
@@
Recent research has explored extending TEEs to accelerators like GPUs, both for x86 (Volos et al., 2018; Jang et al., 2019; Zhu et al., 2020; NVIDIA, 2022a; Lee et al., 2022; Yudha et al., 2022; Ivanov et al., 2023; Mai et al., 2023) and Arm (Jiang et al., 2022; Deng et al., 2022; Wang et al., 2024a) architectures. Most of these approaches involve modifying existing CPU TEEs (especially process-based TEEs like Intel Software Guard Extensions (SGX) and Arm TrustZone (TrustZone)) to secure the GPU software stack, such as GPU drivers, and proposing strategies to ensure the isolation and confidentiality of GPU kernel executions. Apart from this, NVIDIA has introduced a pioneering commercial platform in its Hopper architecture (NVIDIA, 2022a), which supports confidential computing through features like memory encryption and isolated execution contexts. However, the security threats associated with extending TEEs across both CPUs and accelerators in heterogeneous computing systems remain uncertain and require further investigation.
@@
Figure 1. Current state of security research on heterogeneous computing systems with GPU (from the past 15 years). GPU Crypto refers to papers that protect the privacy of GPU-based applications using cryptographic techniques such as HE and MPC. GPU attacks refer to architectural (AA), microarchitectural side-channel/covert-channel (MSCA/MCCA), physical side-channel (PSCA), and software-based fault injection attacks (SFIA).
@@
1999200420092014201920242029 First GPU, NVIDIA GeForce256 ARM TrustZone Launched in ARM11 First Virtualized GPU (NVIDIA VGX) First Attack Exploiting GPU Vuln. (Lee et al., 2014; Maurice et al., 2014) First MSCA on GPU (Luo et al., 2015) Intel SGX Launched in Skylake First MCCA on GPU (Naghibijouybari et al., 2017) AMD SEV Launched in EPYC First FIA on GPU (Frigo et al., 2018) First x86-based GPU TEE (Volos et al., 2018) First MIG-enabled GPU, NVIDIA A100 Intel TDX Announced AMD SEV-SNP Available in Milan First Arm-based GPU TEE (Deng et al., 2022) First Commercial GPU TEE (NVIDIA, 2022a) First MCCA on MIG-enabled GPU (Zhang et al., 2023) First Arm CCA-based GPU TEE (Wang et al., 2024a)
@@ 2.1. General Purpose Graphics Processing Units
Accelerator Year Features Virtualization Shared Memory Hierarchy Memory Size
None L2/L3 Cache DRAM L1 L2 DRAM
@@
Integrated/Discrete GPU. Integrated GPUs, e.g., Intel HD/Iris/UHD/Xe Graphics, Arm, and Qualcomm GPUs in Table 1, are integrated on the same die or package as the CPU. These GPUs share the system’s memory with the CPU. Discrete GPUs are standalone and dedicated graphics processing units separate from the CPU. They have their dedicated memory, which is distinct from the system memory (e.g., NVIDIA and AMD GPUs in Table 1). NVIDIA GPUs have long supported host memory access via PCIe and Unified Memory, but recent architectures like the H100—especially in Grace Hopper configurations—enable tighter CPU-GPU integration with coherent memory sharing over NVLink-C2C (NVIDIA, 2022b).
@@ 2.2. Types of GPU Attacks
2.2.1. Architectural Attacks
These attacks exploit vulnerabilities in the design and configuration of hardware components. This includes misuse of GPU or processor features, privilege escalation, and weak memory isolation. Attackers may target processing units (e.g., processors, hardware accelerators, and co-processors), the memory subsystem, and communication components responsible for interfacing with other system elements to bypass protections or leak sensitive data (Ghasempouri et al., 2023). Logical attacks, a subset of architectural attacks, exploit flaws in the intended logic or functional behavior of the hardware. These may involve undocumented instructions, incorrect permission checks, or logic bugs that allow attackers to alter control flow or data access patterns.
@@ 2.2.4. Fault Injection Attack
2.2.5. Comparison of Attacks on CPUs and GPUs
While attacks on CPU and GPU share some fundamental principles—such as exploiting timing variations, memory leakage, or shared resources—their feasibility, targeted attack surfaces, and required attacker capabilities diverge significantly due to underlying architectural and operational differences. CPUs generally enforce stricter privilege separation and memory isolation, making attacks like residual memory leakage less effective. In contrast, GPUs have historically lacked such safeguards, making them more vulnerable to memory remanence and weak access control. For microarchitectural attacks, GPUs introduce new leakage vectors due to their massive parallelism and SIMT execution model, allowing attackers to exploit fine-grained resource contention (e.g.,, warps, shared caches) at scale. GPUs also expose performance counters and APIs more openly than CPUs, increasing their attack surface. In physical side-channel attacks, GPUs tend to emit more observable signals due to their higher power usage and bandwidth, and their close coupling with CPUs introduces additional leakage through shared interconnects or MMIO. For fault injection, while CPUs are often targeted through physical glitching, GPU attacks frequently exploit software-exposed features like DVFS or power-state manipulation to induce faults during kernel execution. ML workloads on GPUs are resilient to random faults but remain vulnerable to precise, targeted injections. Overall, while CPUs benefit from more mature defense mechanisms, GPU security is still evolving. Adapting proven CPU defenses to GPU-specific designs and threat models remains an important direction for future research.
@@ 2.3.3. Virtualization-based TEEs
Arm CCA. Arm CCA is an advanced isolation technology introduced in Armv9. It preserves TrustZone’s NW and SW while introducing a new realm world designed for running multiple confidential “realms”. These realms, though isolated from other worlds, are managed by untrusted software components (e.g., hypervisor). A lightweight Realm Management Monitor (RMM(Arm, 2023b), operating within the hypervisor layer of the realm world, ensures memory isolation between realms. CCA incorporates a root world to accommodate the highest privilege monitor code stored in firmware (Wang et al., 2024a).
@@ 3. Overview and Taxonomy
Figure 3. Overview of confidential computing on heterogeneous CPU-GPU systems
@@
In Figure 3, we summarise our taxonomy underlying this survey based on relevant related work. In the remainder of this paper, we first review and discuss existing GPU TEEs in Section 4, covering their threat models, hardware modifications, TCB size, and performance overhead. In Section 5, we give an overview of existing attacks targeting GPUs based on the attack types listed in Section 2. We also review state-of-the-art attacks against existing CPU TEEs, such as Intel SGX, Arm TrustZone, and AMD SEV in the supplemental material due to the limited space of this article. Finally, based on these GPU TEE designs and the summarized attacks, we discuss potential attacks against existing (and future) GPU TEEs in Section 6.
@@ 4.1. Trusted Execution Environments on GPUs
TNPU (Lee et al., 2022) Intel x86 NPU Unified Inside CPU enclave SGX Enclave $\checkmark$} $\checkmark$} $\checkmark$ $\checkmark$ $\checkmark$
@@
StrongBox (Deng et al., 2022) Arm Arm Mali GPU Unified Inside normal world TrustZone $\checkmark$} $\checkmark$} $\checkmark$ $\checkmark$
CRONUS (Jiang et al., 2022) Arm General accelerators Dedicated Inside secure world TrustZone virtualization $\checkmark$} $\checkmark$} $\checkmark$ $\checkmark$
@@
Honeycomb (Mai et al., 2023) AMD x86 AMD GPU Dedicated Inside confidential VM AMD SEV-SNP $\checkmark$} $\checkmark$} $\checkmark$ $\checkmark$
CAGE (Wang et al., 2024a) Arm Arm Mali GPU Unified Inside normal world RME, GPC $\checkmark$} $\checkmark$ $\checkmark$
Memory indicates whether the accelerators share the memory with the CPU or not. HW Modification explains if need modification on hardware (● = yes, ○ = no). Secure RM and GPU ME are Secure Resource Management and GPU Memory Encryption. $\dagger$}} indicates that scheme needs additional support (e.g., hardware security features provided in NVIDIA H100 GPU or next-generation Arm devices), as they are not currently considering them.
@@
To support confidential GPU computing on next-generation Arm devices, CAGE (Wang et al., 2024a) presents a new scheme for Arm CCA’s realm-style architecture. It offloads resource-intensive but data-independent tasks to NW, while protecting sensitive data within realms through a shadow task mechanism. By configuring Granule Protection Check (GPC), CAGE restricts access from untrusted components. Additionally, it isolates realms in GPU computing by providing each realm with a distinct GPU memory view in the GPU GPC. CAGE retains high hardware compatibility, as it requires no additional hardware or modifications to existing CPU and GPU configurations.